News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 7 August 2015 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Chandra passed through the tenth and eleventh (and final) eclipses of the season on Jul 31 and Aug 3, respectively, with nominal power and thermal performance. Real-time procedures were executed on Aug 3 to disable SCS-29 and to dump and clear the EPS glitch counters as post-eclipse season actions.

A real-time procedure was executed on Aug 1 to swap the prime solid-state recorder (SSR) from the A-side to the B-side. This swap concludes a six-month interval of health and trending data collection on SSR-A. A real-time procedure was executed on Aug 6 to perform a routine self-check of the Electrical Interface Assembly (EIA) Sequencer.

Of note this week was the announcement on Aug 3 of the release of the Chandra Source Catalog Release 2.0 preliminary detections list. For details see: http://cxc.harvard.edu/csc2/preliminary/

The schedule of targets for the next two weeks is shown below and includes observations of AX J1745.6-2901, which was accepted as a Target of Opportunity on Jul 22, of H1743-322, which was accepted as a Target of Opportunity on Jun 8 and is coordinated with the VLA and NuSTAR, and of the Crab coordinated with HST.



------------------------------------------
CGRaBSJ2334+0736 ACIS-I Aug 9 SPT-CLJ2341-5724 ACIS-I Aug 10 Radiation Belts Aug 11 AXJ1745.6-2901 ACIS-S/HETG SPT-CLJ2341-5724 ACIS-I Aug 13 Radiation Belts WISEPJ190648.47+401 ACIS-S Aug 14 CGCG292-057 ACIS-S Crab ACIS-S Aug 16 L577-71 ACIS-S Radiation Belts H1743-322 ACIS-S/HETG Aug 17 UGC12732 ACIS-S SgrA* ACIS-S PLCKG287+.0+32.9 ACIS-I PSS0955+5940 ACIS-S Aug 18 Radiation Belts Aug 19 Vega ACIS-I Vega ACIS-S SgrAcomplex2 ACIS-I SwiftJ1910.2-0546 ACIS-S Aug 20 RCW38 ACIS-I Aug 21 Radiation Belts NVSSJ225444+134429 ACIS-I Aug 22 Sw1644+57 ACIS-S NGC1892 ACIS-S Aug 23 NGC1688 ACIS-S NGC5044 ACIS-S Radiation Belts Aug 24

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports